### v2

# A characterization of joint completely alternating nets

Abstract.
In this article, we characterize joint completely alternating nets in terms of joint completely monotonic nets. We also discuss completely alternating sequences induced by a class of rational functions. Using the elementary techniques, we obtain a set of conditions sufficient to determine completely alternating sequences interpolated by a class of rational functions. As an application, we obtain a complete characterization of several classes of joint completely monotonic nets induced by rational functions in two variables.
Key words and phrases:
completely monotonic, completely alternating, Hausdorff moment2020 Mathematics Subject Classification:
Primary 44A601. Introduction
Let denote the set of nonnegative real numbers, set of nonnegative integers and set of natural numbers, respectively. For a net of real numbers and for let denotes the forward difference operator which is defined by
where stands for the -tuple with th entry equal to and elsewhere. When we refer as For let denote the operator and denote the sum Here, the operator refers to the identity operator i.e. Given a polynomial represents the degree of A net of real numbers is said to be a joint completely monotonic if
and joint completely alternating if
For we call the sequence as completely monotonic sequence and completely alternating sequence if it satisfies the first and second equation, respectively. Below, we enlist well-known characterizations of completely monotonic and completely alternating sequence in terms of measure (see [11, Propositions 6.11, 6.12]).
Theorem 1.1.
Let be a positive integer and be a net of real numbers.
The net is joint completely monotonic if and only if it is a Hausdorff moment net, that is, there exists a positive Borel measure on such that
(1.1) -
The net is joint completely alternating if and only if there exist and a positive Borel measure on such that
where represents the standard inner product on
Note that a joint completely alternating net is invariant under a shift by a constant. Due to this fact, in this article, we work with joint completely alternating nets of positive real numbers.
Recall that the measure in (1.1) is called the representing measure of and it is unique. This is a consequence of the -dimensional Weierstrass theorem and the Riesz representation theorem. In [13], F. Hausdorff characterized a completely monotonic sequence by Hausdorff moment sequence and the above characterization about completely alternating sequence was first obtained as a special case of the Levy-Khinchin representation theory on abelian semigroups (see [9], [15]). For a comprehensive study on these topics, we refer the reader to [9], [11], [15], [16].
A lot of work has been done on the characterization of completely monotonic sequences in terms of Hausdorff moment sequences, for instance, [10, 12, 14] and references therein. However, the characterization of a completely monotonic sequence induced by rational functions in terms of zeroes and poles is challenging and seems to be beyond the limit of present understanding. One of the important results in this direction is the sufficient conditions to identify completely monotonic sequences introduced by Ball [7] which are solely dependent on the zeroes and poles of the rational functions. This article aims to determine a certain completely alternating sequences interpolated by rational functions with real zeros and poles in terms of sufficient conditions. In this article, we deal with the following class of rational functions, that is, for
where is a polynomial of degree at most , and ’s are distinct positive real numbers arranged in non-decreasing order.
In this article, more precisely, we concentrate on the following moment problem
Question 1.2.
Among all the sequences of rational functions , defined in (1.3), which are completely alternating?
From the operator theoretic aspect, completely monotonic and completely alternating sequences are used in revealing the connection with subnormal and completely hyperexpansive operators, respectively (see [1], [2], [3], [4], [5], [6], [8]).
Before proceeding further, we encounter a non-example which shows that any sequence of rational functions is not completely alternating sequence.
Example 1.3.
Let Note that
Since is not a completely alternating sequence.
The above example motivates us to identify the completely alternating sequence among the sequence of rational functions of the form (1.3).
We now state the main results of this paper. The first one characterizes joint completely alternating nets and the second result involves a set of conditions sufficient to determine the completely alternating sequences induced by rational functions.
Theorem 1.4.
Let be a net of positive real numbers. Then the following are equivalent
the net is joint completely alternating,
is a joint completely monotonic net.
Theorem 1.5.
Let and be a strictly increasing sequence of positive real numbers. Consider the polynomial given by Let be any polynomial of degree at most with the coefficient of non-negative. Then, the sequence is completely alternating provided
where
As an application of this result, we also develop a set of conditions sufficient to identify the completely alternating sequences induced by rational functions defined in (1.3), with and or , which solely depend on its zeros and poles (see Corollary 3.4, also see Theorem 3.1). These sufficient conditions are very easy to check as well as turn out to be necessary in a sense discussed in Remark 3.5. In a recent article [4], the authors have studied joint complete monotonicity of the net (see [4, Theorem 1.4]) and solved the Cauchy dual subnormality problem for a class of toral -isometric weighted -shifts which are separate -isometries. Combining Theorem 1.4 and 3.1, we also obtain a complete characterization of several classes of joint completely monotonic nets induced by rational functions in two variables (see Corollary 3.6).
2. A proof the Theorem 1.4
In this section, we present a proof of Theorem 1.4. In [5, Proposition 6] (also see [12, Theorem 2.5]), the authors have shown a correspondence between completely alternating sequences and completely monotonic sequences. It is worth mentioning here that there are only a few interesting examples of joint complete monotonic nets in more than one variable in the literature. Our result in this section allow us to construct non-trivial examples from the joint completely alternating nets (see Corollary 3.6).
For the reader’s convenience, we recall the following result which we need in the proof of Theorem 1.4 ([11, Proposition 6.10, p. 133]).
Proposition 2.1.
Let Then the net is joint completely alternating if and only if is a joint completely monotonic net for all
We are now ready to prove Theorem 1.4.
Proof of Theorem 1.4.
To prove the equivalence, we note the following identity:
(i)(ii): Assume that is a joint completely alternating sequence. By Proposition 2.1, is a joint completely monotonic net, for every Equivalently, the net is joint completely monotonic for every This together with the fact that is a completely monotonic sequence yields that for every
Combining this with (2.1), we get
This shows that is a joint completely monotonic net.
(ii)(i): Assume that is a joint completely monotonic net. This together with (2.1) yields that for every
By the joint complete monotononicity of the net it is easy to see that the sequence
is completely monotonic for every Therefore its representing measure which is a weighted Lebesgue measure with weight function is positive for almost every But since is continuous on we have
This yields that for every is a joint completely monotonic net and equivalently, the net is joint completely alternating. ∎
3. completely alternating sequences induced by rational functions
In this section, we present a proof of Theorem 1.5. Along with that, we show that every member of a class of sequences induced by rational functions is completely alternating if it satisfies a set of conditions developed solely in terms of its zeros and poles.
Note that if the measure of the completely alternating sequence in (1.2) is finite then (1.2) can be rewritten as
for some and We will make use of this expression throughout this section.
We begin with the following result, which deals with sequences generated by rational functions of at most degree 2 and finds out when it will be a completely alternating sequence in terms of necessary and sufficient conditions. It has a deep impact on identifying a class of completely alternating sequences.
Theorem 3.1.
Let with and , then we have the following:
Let Then the sequence is completely alternating if and only if
Let Then the sequence is completely alternating if and only if
Proof.
We divide the proof into two cases. Assume that By the partial fraction decomposition,
where and Note that
In view of Theorem 1.1(ii), it suffices to check that for all for the proof of both the necessity and sufficiency part. By [4, Lemma 3.4], this is equivalent to and Further, it is equivalent to Now if then again by the partial fraction decomposition,
where and Note that
It is easy to verify that if and only if and Latter conditions are equivalent to and This completes the proof.
Analogous argument, as used in the first part, works here as well.
The following lemma plays a crucial role in proving Theorem 1.5.
Lemma 3.2.
Let For any collection of real numbers satisfying and any collection of strictly increasing positive real numbers, the function
Proof.
We use the induction on to prove the lemma. For since Assume the induction hypothesis i.e. the statement of the lemma is true for any Now consider where and are as in the statement. Let If is empty, then we are done. Let be the largest element of Let if possible Then For every
By the induction hypothesis, we have
If Then We divide this case into the following two cases.
Case (i):
Note that for every
The last inequality follows from the induction hypothesis.
Case (ii):
Note that for every
The last inequality follows from the induction hypothesis. This completes the proof. ∎
Next, we present the proof of Theorem 1.5 by generalizing the idea used in the proof of Theorem 3.1.
Proof of Theorem 1.5.
Note that by partial fraction decomposition, we have
where and is the coefficient of in which is non negative. It is easy to see using (3.1) that
where To prove the sequence is completely alternating, by Theorem 1.1(ii), it is now sufficient to check that for This follows from Lemma 3.2. This completes the proof. ∎
Remark 3.3.
In Theorem 1.5, turns out to be a necessary condition. Otherwise, if the sequence corresponding to the rational function with is completely alternating then the sequence turns out to be divergent, that is,
is not finite. But on the other hand, according to the Theorem 1.1(ii), for a completely alternating sequence , where is a non negative real number as in (1.2). This also explains that the non-negativity of the coefficient of in Theorem 1.5 is also necessary.
As an application to Theorem 1.5, we obtain the following classes of completely alternating sequences.
Corollary 3.4.
Let and be any of the following
with
with
Then the sequence is completely alternating.
Proof.
(i) Note that
where This can be rewritten as
(we assume that the product over the empty set is ). Under the given conditions, it is easy to see that By Theorem 1.5, the sequence is completely alternating. A similar proof works for (ii). ∎
Remark 3.5.
The sufficient conditions in Corollary 3.4 turn out to be necessary if we further assume that, in the first case where the degree of is same as the degree of , the sequences for each and are complete alternating and in the second case where the degree of is bigger than the degree of by 1, the sequences for each are completely alternating. The proof follows from the Theorem 3.1.
The following corollary presents a complete characterization of certain classes of joint completely monotonic nets in two variables. This is an application of Theorems 1.4, 3.1 (also see [4][Theorems 3.1, 3.3, Question 1.1]).
Corollary 3.6.
Let Then the following statements are valid
is a joint completely monotonic net if and only if
The net is joint completely monotonic if and only if
is a joint completely monotonic net if and only if
Acknowledgements: The authors are grateful to Prof. Sameer Chavan for some insightful comments/suggestions. The second named author acknowledges the Department of Mathematics, Birla Institute of Technology and Science K.K. Birla Goa Campus for warm hospitality and is grateful to the DST-INSPIRE Faculty Fellowship. The first-named author’s research is supported by the DST-INSPIRE Faculty Fellowship No. DST/INSPIRE/04/2020/001250.

### v3

@@
-A characterization of joint completely alternating nets
+A characterization of completely alternating functions
 Abstract.
-In this article, we characterize joint completely alternating nets in terms of joint completely monotonic nets.
-We also discuss completely alternating sequences induced by a class of rational functions.
-Using the elementary techniques, we obtain a set of conditions sufficient to determine completely alternating sequences interpolated by a class of rational functions.
-As an application, we obtain a complete characterization of several classes of joint completely monotonic nets induced by rational functions in two variables.
+In this article, we characterize completely alternating functions on an abelian semigroup in terms of completely monotone functions on the product semigroup .
+We also discuss completely alternating sequences induced by a class of rational functions and obtain a set of sufficient conditions (in terms of it’s zeros and poles) to determine them.
+As an application, we show a complete characterization of several classes of completely monotone functions on induced by rational functions in two variables.
+We also derive a set of necessary conditions for the complete monotonicity of the sequence .
 Key words and phrases:
-completely monotonic, completely alternating, Hausdorff moment2020 Mathematics Subject Classification:
+completely monotone, completely alternating, Hausdorff moment2020 Mathematics Subject Classification:
 Primary 44A601. Introduction
-Let denote the set of nonnegative real numbers, set of nonnegative integers and set of natural numbers, respectively.
-For a net of real numbers and for let denotes the forward difference operator which is defined by
-where stands for the -tuple with th entry equal to and elsewhere.
-When we refer as For let denote the operator and denote the sum Here, the operator refers to the identity operator i.e.
-Given a polynomial represents the degree of A net of real numbers is said to be a joint completely monotonic if
-and joint completely alternating if
-For we call the sequence as completely monotonic sequence and completely alternating sequence if it satisfies the first and second equation, respectively.
-Below, we enlist well-known characterizations of completely monotonic and completely alternating sequence in terms of measure (see [11, Propositions 6.11, 6.12]).
+Let be an abelian semigroup with identical involution and denotes the set of real-valued functions on For define the shift operator on by and the backward difference operator where denotes the identity operator on A function is called completely monotone if it is nonnegative and if for all finite set and
+A function is called completely alternating if for all finite set and
+The set of completely monotone and completely alternating functions on is denoted by and , respectively.
+Given a set let denote the Cartesian product of with itself times.
+Let and denote the set of non-negative integers and the set of non-negative real numbers, respectively.
+Below, we enlist well-known characterizations of completely monotone and completely alternating functions on the abelian semigroup , in terms of measure (see [12, Propositions 6.11, 6.12, p.
+134]).
 Theorem 1.1.
-Let be a positive integer and be a net of real numbers.
-The net is joint completely monotonic if and only if it is a Hausdorff moment net, that is, there exists a positive Borel measure on such that
+For the following holds:
+if and only if it is a Hausdorff moment net, that is, there exists a positive Borel measure on such that
 (1.1) -
-The net is joint completely alternating if and only if there exist and a positive Borel measure on such that
+if and only if there exist and a positive Borel measure on such that
 where represents the standard inner product on
-Note that a joint completely alternating net is invariant under a shift by a constant. Due to this fact, in this article, we work with joint completely alternating nets of positive real numbers.
+Note that a completely alternating function is invariant under a shift by a constant. Due to this fact, in this article, we work with completely alternating functions which are positive.
 Recall that the measure in (1.1) is called the representing measure of and it is unique.
@@
 In [14], F.
-Hausdorff characterized a completely monotonic sequence by Hausdorff moment sequence and the above characterization about completely alternating sequence was first obtained as a special case of the Levy-Khinchin representation theory on abelian semigroups (see [9], [15]).
+Hausdorff characterized a completely monotone sequence by Hausdorff moment sequence and the above characterization about completely alternating sequence was first obtained as a special case of the Levy-Khinchin representation theory on abelian semigroups (see [10], [16]).
 For a comprehensive study on these topics, we refer the reader to [10], [12], [16], [17].
-A lot of work has been done on the characterization of completely monotonic sequences in terms of Hausdorff moment sequences, for instance, [10, 12, 14] and references therein.
-However, the characterization of a completely monotonic sequence induced by rational functions in terms of zeroes and poles is challenging and seems to be beyond the limit of present understanding.
-One of the important results in this direction is the sufficient conditions to identify completely monotonic sequences introduced by Ball [7] which are solely dependent on the zeroes and poles of the rational functions.
-This article aims to determine a certain completely alternating sequences interpolated by rational functions with real zeros and poles in terms of sufficient conditions.
+A lot of work has been done on the characterization of completely monotone functions in terms of Hausdorff moment sequences, for instance, [11, 13, 15] and references therein.
+However, the characterization of a completely monotone sequence induced by rational functions in terms of zeroes and poles is challenging and seems to be beyond the limit of present understanding.
+One of the important results in this direction is the sufficient conditions to identify completely monotone sequences introduced by Ball [8] which are solely dependent on the zeroes and poles of the rational functions (see Theorem 1.6).
+This article aims to determine certain completely alternating sequences interpolated by rational functions with real zeros and poles in terms of sufficient conditions.
 In this article, we deal with the following class of rational functions, that is, for
-where is a polynomial of degree at most , and ’s are distinct positive real numbers arranged in non-decreasing order.
-In this article, more precisely, we concentrate on the following moment problem
+where is a polynomial of degree at most and
+Here, more precisely, we concentrate on the following moment problem
 Question 1.2.
 Among all the sequences of rational functions , defined in (1.3), which are completely alternating?
-From the operator theoretic aspect, completely monotonic and completely alternating sequences are used in revealing the connection with subnormal and completely hyperexpansive operators, respectively (see [1], [2], [3], [4], [5], [6], [8]).
+From the operator theoretic aspect, completely monotone and completely alternating sequences are used in revealing the connection with subnormal and completely hyperexpansive operators, respectively (see [1], [2], [3], [4], [7], [6], [9]).
 Before proceeding further, we encounter a non-example which shows that any sequence of rational functions is not completely alternating sequence.
 Example 1.3.
-Let Note that
+Let For the semigroup we write Note that
 Since is not a completely alternating sequence.
-The above example motivates us to identify the completely alternating sequence among the sequence of rational functions of the form (1.3).
+The above example motivates us to identify completely alternating sequences among the sequence of rational functions of the form (1.3).
 We now state the main results of this paper.
-The first one characterizes joint completely alternating nets and the second result involves a set of conditions sufficient to determine the completely alternating sequences induced by rational functions.
+The first one characterizes a completely alternating function on an abelian semigroup with the identical involution, and the second result involves a set of conditions sufficient to determine the completely alternating sequences induced by rational functions.
 Theorem 1.4.
-Let be a net of positive real numbers. Then the following are equivalent
-the net is joint completely alternating,
-is a joint completely monotonic net.
+Let S be an abelian semigroup with the identical involution and let be a strictly positive function. The function defined by
+is completely monotone on the product semigroup with the identical involution if and only if is completely alternating on i.e.
 Theorem 1.5.
@@
 As an application of this result, we also develop a set of conditions sufficient to identify the completely alternating sequences induced by rational functions defined in (1.3), with and or , which solely depend on its zeros and poles (see Corollary 3.4, also see Theorem 3.1).
-These sufficient conditions are very easy to check as well as turn out to be necessary in a sense discussed in Remark 3.5.
-In a recent article [4], the authors have studied joint complete monotonicity of the net (see [4, Theorem 1.4]) and solved the Cauchy dual subnormality problem for a class of toral -isometric weighted -shifts which are separate -isometries.
-Combining Theorem 1.4 and 3.1, we also obtain a complete characterization of several classes of joint completely monotonic nets induced by rational functions in two variables (see Corollary 3.6).
+These sufficient conditions are very easy to check.
+They turn out to be necessary in a sense discussed in Remark 3.5.
+In a recent article [4], the authors have studied complete monotonicity of the net (see [4, Theorem 1.4]) and solved the Cauchy dual subnormality problem for a class of toral -isometric weighted -shifts which are separate -isometries.
+Combining Theorem 1.4 and 3.1, we also obtain a complete characterization of several classes of completely monotone nets induced by rational functions in two variables (see Corollary 3.9).
+Also, we recover Theorem 3.1 in [4] with another characterization of completely monotone sequences as an application of Theorem 1.4 (see Corollary 3.10).
+Next we turn our attention to the complete monotonicity of the following class of rational function
+where and In [8], K. Ball obtain a set of sufficient conditions in terms of zeros and poles of (1.4) for which (1.4) is completely monotone. For the reader’s convenience, we recall the result here:
+Theorem 1.6 (K. Ball, 1994).
+To the best of our knowledge, it was not known whether the Ball’s sufficiency conditions for the complete monotonicity of (1.4) are necessary.
+A possible reason could be that these conditions are both necessary and sufficient for and (to be seen in Section 4).
+It was mentioned in his paper, that those conditions (1.5) are almost necessary (see [8, Page 3]) that is, if each is completely monotone then (1.5) hold.
+This follows from [8, Remark].
+In Section 4, we prove that Ball’s sufficiency conditions are not necessary and obtain a new set of necessary conditions for this class which is a generatization of Ball’s sufficiency condition.
+Before we state this result, we need the following notation.
+Let denote the set of all permutations on the set .
+Theorem 1.7.
+Let and be two non-decreasing sequences of positive real numbers. Let and Then
+If the sequence is completely monotone then there exists a with such that
+(1.6) for every
+If the sequence is completely alternating then there exists a with such that
+for every
 2. A proof the Theorem 1.4
 In this section, we present a proof of Theorem 1.4.
-In [5, Proposition 6] (also see [12, Theorem 2.5]), the authors have shown a correspondence between completely alternating sequences and completely monotonic sequences.
-It is worth mentioning here that there are only a few interesting examples of joint complete monotonic nets in more than one variable in the literature.
-Our result in this section allow us to construct non-trivial examples from the joint completely alternating nets (see Corollary 3.6).
-For the reader’s convenience, we recall the following result which we need in the proof of Theorem 1.4 ([11, Proposition 6.10, p. 133]).
+In [7, Proposition 6] (also see [13, Theorem 2.5]), the authors have shown a correspondence between completely alternating sequences and completely monotone sequences.
+As far as we know, there are only a few interesting examples of complete monotone functions in more than one variable in the literature.
+Our result in this section allows us to construct non-trivial examples from completely alternating functions (see Corollary 3.9).
+For the reader’s convenience, we recall the following result which we need in the proof of Theorem 1.4 ([12, Proposition 6.10, p. 133]).
 Proposition 2.1.
-Let Then the net is joint completely alternating if and only if is a joint completely monotonic net for all
+Let Then if and only if for all
 We are now ready to prove Theorem 1.4.
@@
 To prove the equivalence, we note the following identity:
-(i)(ii): Assume that is a joint completely alternating sequence.
-By Proposition 2.1, is a joint completely monotonic net, for every Equivalently, the net is joint completely monotonic for every This together with the fact that is a completely monotonic sequence yields that for every
-Combining this with (2.1), we get
-This shows that is a joint completely monotonic net.
-(ii)(i): Assume that is a joint completely monotonic net. This together with (2.1) yields that for every
-By the joint complete monotononicity of the net it is easy to see that the sequence
-is completely monotonic for every Therefore its representing measure which is a weighted Lebesgue measure with weight function is positive for almost every But since is continuous on we have
-This yields that for every is a joint completely monotonic net and equivalently, the net is joint completely alternating. ∎
+: Assume that . By Proposition 2.1, for every Equivalently, for every This together with the fact that is completely monotone yields that for every
+: Assume that This together with (2.1) yields that for every
+Since
+for every finite set Note that for a finite set the function This can be seen by taking in (2.2).
+Therefore the representing measure of the sequence is a weighted Lebesgue measure (by (2.2)) with weight function is positive for almost every But since is continuous on we have
+This yields that for every and equivalently, ∎
+The following result is a consequence of Theorem 1.4.
+Corollary 2.2.
+Assume the notations of Theorem 1.4 and Then if and only if
+Proof.
+This follows from Theorem 1.4 and the fact that multiplying by a positive real number does not change the completely alternating (resp. completely monotone) property of a function. ∎
 3. completely alternating sequences induced by rational functions
@@
 Note that if the measure of the completely alternating sequence in (1.2) is finite then (1.2) can be rewritten as
-for some and We will make use of this expression throughout this section.
+for some and Conversely, if where and are polynomials with and then can be written as in (3.1) (using partial fraction decomposition) with being a finite signed measure (since is finite).
+For irreducible with zeros in the left half-plane, one can use the method from [1, p.
+800] to derive an expression like (3.1).
+We will make use of this expression throughout this section.
 We begin with the following result, which deals with sequences generated by rational functions of at most degree 2 and finds out when it will be a completely alternating sequence in terms of necessary and sufficient conditions.
@@
 Let Then the sequence is completely alternating if and only if
+Let Then the sequence is completely alternating if and only if
 Proof.
+This can be seen from the following identity:
 We divide the proof into two cases. Assume that By the partial fraction decomposition,
-where and Note that
-In view of Theorem 1.1(ii), it suffices to check that for all for the proof of both the necessity and sufficiency part.
-By [4, Lemma 3.4], this is equivalent to and Further, it is equivalent to Now if then again by the partial fraction decomposition,
+where and we note that
+Since is finite, the function is Lebesgue integrable on In view of Theorem 1.1(ii), it suffices to check that for all to establish both necessity and sufficiency.
+This condition is equivalent to requiring and Further, it is equivalent to Now if then again by the partial fraction decomposition,
 where and Note that
 It is easy to verify that if and only if and Latter conditions are equivalent to and This completes the proof.
-Analogous argument, as used in the first part, works here as well.
+Analogous argument, as used in the second part, works here as well.
 The following lemma plays a crucial role in proving Theorem 1.5.
@@
 where and is the coefficient of in which is non negative. It is easy to see using (3.2) that
-where To prove the sequence is completely alternating, by Theorem 1.1(ii), it is now sufficient to check that for This follows from Lemma 3.2. This completes the proof. ∎
+where To prove the sequence is completely alternating, by (3.1), it is now sufficient to check that for This follows from Lemma 3.2. This completes the proof. ∎
 Remark 3.3.
@@
 This also explains that the non-negativity of the coefficient of in Theorem 1.5 is also necessary.
-As an application to Theorem 1.5, we obtain the following classes of completely alternating sequences.
+As an application of Theorem 1.5, we obtain the following classes of completely alternating sequences.
 Corollary 3.4.
@@
 The proof follows from the Theorem 3.1.
-The following corollary presents a complete characterization of certain classes of joint completely monotonic nets in two variables.
+The next result helps in finding more classes of completely alternating sequences where not all ’s are negative.
+Proposition 3.6.
+For any rational function let denote the coefficient in the partial fraction decomposition of Then
+Proof.
+We use induction to prove the statement. For
+So the statement is true for since Assume that the statement is true for any Then
+From this expression, we note that From the induction hypothesis, we obtain ∎
+Now we present some classes of completely alternating sequences for which not all ’s are negative.
+Corollary 3.7.
+Let and with Then the sequence is completely alternating.
+Proof.
+Remark 3.8.
+The following corollary presents a complete characterization of certain classes of completely monotone nets in two variables.
 This is an application of Theorems 1.4, 3.1 (also see [4][Theorems 3.1, 3.3, Question 1.1]).
@@
 Let Then the following statements are valid
-is a joint completely monotonic net if and only if
-The net is joint completely monotonic if and only if
-is a joint completely monotonic net if and only if
+if and only if
+if and only if
+if and only if
+As an immediate application of Theorem 1.4, we are able to recover Theorem 3.1 in [4] and along with that, we obtain another equivalent statement.
+Corollary 3.10.
+For the following are equivalent:
+The net is completely monotone.
+The net is completely monotone.
+4. Necessary conditions
+In this section, we study the necessary conditions for the complete monotonicity of (1.4).
+As a consequence, we also obtain necessary conditions for the sequences of the type (1.4) which are completely alternating.
+We first show that Ball’s sufficiency conditions for the complete monotonicity of (1.4) are not necessary and then present a proof of Theorem 1.7 with the necessary results required to prove.
+Example:
+Let and For every
+Let Note that
+This yields that is decreasing on Since This, together with (4.1) shows that is a completely monotone sequence.
+Equivalently, the function is completely monotone (see [5, Proposition 6]).
+But this does not satisfy Ball’s sufficiency conditions.
+The following lemma presents a simple characterization of the complete monotonicity of a special class.
+A proof in a more general setting of the following lemma has been obtained in [15, Theorem 4.1].
+We present an elementary proof of the following lemma, and this will play an essential role in proving Theorem 1.7.
+Lemma 4.1.
+Let and Let and Then, is completely monotone if and only if
+Proof.
+For a proof of the sufficiency part, assume that It is easy to see that, is a sum of two completely monotone sequence, hence is completely monotone.
+Since the product of completely monotone sequences is completely monotone, the sequence is completely monotone.
+For a proof of necessity part, assume that the sequence is completely monotone.
+By multiplying, for any if necessary, we can assume that all ’s are distinct.
+Let, if possible, For every
+By the partial fraction decomposition, there exists such that
+Observe that Note that the representing measure of is the weighted Lebesgue measure with weight function given by
+It is easy to see that there exists such that This contradicts the assumption. Hence ∎
+We recall the following result which will be crucial in proving Theorem 1.7.
+Proposition 4.2.
+Let and be a non-decreasing sequence of positive real numbers. Let Assume that is completely monotone. Then,
+Proof.
+See [8, Remark]. ∎
+Now, K. Ball’s sufficiency conditions combined with Propositions 4.2 and Lemma 4.1 provide a complete characterization of completely monotone rational functions of the type (1.4) in the case of
+Proposition 4.3.
+Let be polynomials given by and with . Then, is completely monotone if and only if
+We are now ready to prove Theorem 1.7.
+Proof of Theorem 1.7.
+(i) We use the finite induction on .
+For the base cases, take and follows from Proposition 4.3.
+Now, assume the induction hypothesis.
+Let if possible there exists a such that (otherwise the result is true trivially).
+Clearly by Lemma 4.1.
+Also note that, the product of two completely monotone sequences is completely monotone.
+Now, by using these facts, we conclude that the sequence
+is completely monotone. Now take and Similarly take and By induction hypothesis there exists such that and
+Let be such that Then by (4.2), we have
+By taking we have which is the desired permutation. By Proposition 4.2 and (4.3), we obtain (1.6).
+(ii) This follows from (i) and the fact that if the sequence of positive real numbers is completely alternating then is completely monotone.
+It might be tempting to believe that the new necessary conditions for the complete monotonicity of (1.4) are sufficient.
+But we now present an example which shows that the necessary conditions obtained in Theorem 1.7 are not sufficient.
+Example: Let and For every
+Let It is easy to verify that for all This together with (4.4) yields that is not completely monotone.
 Acknowledgements: The authors are grateful to Prof.